# Learning Physical Systems: Symplectification via Gauge Fixing in Dirac Structures

### v1

@@ I Introduction
$T^{*}\widetilde{Q}:=T^{*}Q\times T^{*}\!\bigl{(}\mathbb{R}\times\mathbb{R}^{m}\bigr{)},$ (1)
$(\mathbf{Q},\mathbf{P}):=(q^{0},q^{i},\lambda_{a};\,p_{0},p_{i},\pi_{a}),$
@@
which is closed and non-degenerate by construction. Choosing the extended Hamiltonian $\widetilde{H}(\mathbf{Q},\mathbf{P})=H(\mathbf{q},\mathbf{p})+p_{0}+\lambda_{a}\,\phi^{a}(\mathbf{q})$, where $\frac{d\pi_{a}}{dt}=-\phi^{a}(\mathbf{q})=0$ encodes the original constraints, Hamilton’s equations on $(T^{*}\widetilde{Q},\widetilde{\Omega})$ reproduce the constrained dynamics once the Dirac gauge:
@@ II-A Problem Statement
where $\mathbf{M}(\mathbf{q})$ represents the mass matrix, $\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})$ the nonlinear dynamics terms, $\mathbf{B}$ the control-input ($\mathbf{u})$ selection matrix, $\mathbf{J}_{c}(\mathbf{q})$ the concatenated contact Jacobians and $\mathbf{F}_{c}$ the contact wrenches. The scleronomic contact constraint $\phi_{i}(\mathbf{q})=0$ is active only for the feet ($i$) that are in contact, within the contact set $\mathcal{I}_{c}$. The configuration space $(Q)$ comprises a composite lie group $\mathbb{SE}(3)\times\mathbb{R}^{12}$, while its tangent space $TQ$ lives in the euclidean $\mathbb{R}^{18}$ space. After the Dirac lift, the system is characterized by a lifted 87-dimensional augmented phase-space.
@@ II-C Flow-Matching Objective
$\mathcal{L}_{\text{FM}}=\frac{1}{|\mathcal{D}|}\sum_{t}\left|(\mathbf{v}^{*}-\Psi_{\theta}(\hat{\mathbf{z}}_{t})\right|_{2}^{2},$ (9)
which reduces to, and generalizes, the exact transport equation used by [20].
@@ II-D SympNet Step Predictor
$\mathbf{z}_{t+\Delta}=\mathbf{z}_{t}\oplus\Delta S_{\phi}\bigl{(}\mathbf{z}_{t}\bigr{)},\qquad S_{\phi}^{*}\widetilde{\Omega}=\widetilde{\Omega}.$ (10)
@@
$\mathcal{L}_{\text{pred}}=\frac{1}{|\mathcal{D}|}\sum_{t}\left|\mathbf{z}_{t+\Delta}\ominus\mathbf{x}_{t+\Delta}\right|_{2}^{2},$ (11)
while the PSN parameters are frozen, guaranteeing that the learned lift remains unchanged. The $\oplus$ and $\ominus$ indicate the phase-space’s integrate and difference operators, since the state of the floating-base quadruped constitutes a composite Lie-Group.

### v2

@@ I Introduction
$T^{*}\widetilde{Q}=T^{*}Q\times T^{*}\!\bigl(\mathbb{R}\times\mathbb{R}^{m}\bigr),$ (1)
$(\mathbf{Q},\mathbf{P})=(q^{0},q^{i},\lambda_{a};\,p_{0},p_{i},\pi_{a}),$
@@
[…]
@@ II-A Problem Statement
where $\mathbf{M}(\mathbf{q})$ represents the mass matrix, $\mathbf{C}(\mathbf{q},\dot{\mathbf{q}})$ the nonlinear dynamics terms, $\mathbf{B}$ the control-input ($\mathbf{u})$ selection matrix, $\mathbf{J}_{c}(\mathbf{q})$ the concatenated contact Jacobians and $\mathbf{F}_{c}$ the contact wrenches. The scleronomic contact constraint $\phi_{i}(\mathbf{q})=0$ is active only for the feet ($i$) that are in contact, within the contact set $\mathcal{I}_{c}$. The configuration space $(Q)$ comprises a composite lie group $\mathbb{SE}(3)\times\mathbb{R}^{12}$, while its tangent space $TQ$ lives in the euclidean $\mathbb{R}^{18}$ space. After the Dirac lift, the system is characterized by a lifted 86-dimensional augmented phase-space with the base orientation represented as quaternion.
@@ II-C Flow-Matching Objective
$\mathcal{L}_{\text{FM}}=\frac{1}{|\mathcal{D}|}\sum_{t}\lVert\mathbf{v}^{*}-\Pi(\mathbf{v}_{t})\rVert_{2}^{2},$ (9)
where $\Pi$ is a projection operator that maps the full augmented velocity $\mathbf{v}_{t}$ to the physical phase space coordinates, which reduces to, and generalizes, the exact transport equation used by [20].
@@ II-D SympNet Step Predictor
$\mathbf{z}_{t+\Delta}=S_{\phi}\bigl(\mathbf{z}_{t}\bigr),\qquad S_{\phi}^{*}\widetilde{\Omega}=\widetilde{\Omega}$. (10)
@@
$\mathcal{L}_{\text{pred}}=\frac{1}{|\mathcal{D}|}\sum_{t}\lVert\Pi(\mathbf{z}_{t+\Delta})\ominus\mathbf{x}_{t+\Delta}\rVert_{2}^{2},$ (11)
while the PSN parameters are frozen, guaranteeing that the learned lift remains unchanged. The $\ominus$ operator computes the difference between the projected predicted state $\Pi(\mathbf{z}_{t+\Delta})$ and the ground-truth physical state $\mathbf{x}_{t+\Delta}$, appropriately handling the composite Lie-Group of the floating-base quadruped.